OpenStudy (anonymous):

Graph the system of constraints and find the value of x and y that maximize the objective function. x>0 y>0 y<1/5x+2 5>y+x

OpenStudy (raden):

can u graph the lines of y=1/5x+2 and 5=y+x

OpenStudy (anonymous):

it always confuses me